# Categories of Line Defects and Cohomological Hall Algebras

### v2

@@ 5.1 Framed quivers and BPS degeneracies
This is an extension! In other words, $A$ is an ${\mathcal{O}}^{*}$-equivariant submodule of the tensor product, but $\mathfrak{q}^{-1}X_{\gamma_{1}}A$ is not and we get a non-trivial exact sequence
@@ 6.2 Simple products
The auxiliary bimodule map $\pi_{\vec{f}_{a},\vec{f}_{a+1}}$ that we defined above indeed maps $b\otimes b\to b$. Then we still need to show that $\pi_{\vec{f}_{a},\vec{f}_{a+1}}$ actually lands in $\mathrm{sH}(Q;\vec{f}_{a}+\vec{f}_{a+1})$ rather than $\mathrm{Sh}(Q;\vec{f}_{a}+\vec{f}_{a+1})$, i.e. that we do not encounter images of the form $bab$ where $a$ is a non-trivial algebra element that cannot be brought to the left and/or to the right of the $b$’s.
@@ 6.3 Interesting tensor products
    $[e^{(1)}_{n}\,,\,b_{\pm\gamma_{1}}]=0\qquad\textrm{and}\qquad e^{(2)}_{n}\,b_{\pm\gamma_{1}}=b_{\pm\gamma_{1}}\,e^{(2)}_{n\mp 1}\,.$ (239)
@@ 6.3.3 $B_{2}$ and $B_{4}$
anti-commutes with all generators in $A$ [^20] If we had done this computation directly in terms of modes using only the exchange relations, we would have $\{e^{(1)}_{n\geq 1},u_{24}\}=\{e^{(2)}_{n},u_{24}\}=0$, and $\{e^{(1)}_{0},u_{24}\}=-b_{\gamma_{1}}\otimes c_{0}$, which vanishes only upon invoking the cubic relation (263). Therefore we get the expected
@@ 6.3.5 $B_{4}$ and $B_{1}$
anti-commutes with all $e^{(1)}_{n}$’s and commutes with all $e^{(2)}_{n}$’s. [^21] Similar to the $B_{2}\otimes_{A}B_{4}$ case, if we had done this computation directly in terms of modes using only the exchange relations, we would have $\{e^{(1)}_{n},u_{41}\}=[e^{(2)}_{n\geq 1},u_{41}]=0$, and $[e^{(2)}_{0},u_{41}]=-c_{0}\otimes b_{\gamma_{2}}$, which vanishes only upon invoking the cubic relation (263). Therefore we get the exact sequences
@@ 7.1 Mutation orbits
$[-1,0,1]\quad\rightarrow\quad[1,-1,0]\quad\rightarrow\quad[0,1,-1]\,.$ (285)
@@
$\vec{f}_{0}^{+}=[1,0,0]\,,\qquad\quad$ $\vec{f}_{0}^{0}=[-1,0,1]\,,\qquad\quad\vec{f}_{0}^{-}=[0,0,-1]\,,$ (286)
$\vec{f}_{1}^{+}=[0,1,0]\,,\qquad\quad$ $\vec{f}_{1}^{0}=[1,-1,0]\,,\qquad\quad\vec{f}_{1}^{-}=[-1,0,0]\,,$ (287)
$\vec{f}_{2}^{+}=[0,0,1]\,,\qquad\quad$ $\vec{f}_{2}^{0}=[0,1,-1]\,,\qquad\quad\vec{f}_{2}^{-}=[0,-1,0]\,,$ (288)
@@
$\vec{f}_{0}^{+}=[1,0,0]\,,\qquad\quad$ $\vec{f}_{0}^{0}=[0,-1,0]\,,\,\,\,\,\qquad\,\,\,\vec{f}_{0}^{-}=[0,1,-1]\,,$ (289)
$\vec{f}_{1}^{+}=[0,0,1]\,,\qquad\quad$ $\vec{f}_{1}^{0}=[1,0,-1]\,,\qquad\quad\,\vec{f}_{1}^{-}=[-1,0,0]\,,$ (290)
$\vec{f}_{2}^{+}=[1,-1,0]\,,\qquad\quad\,$ $\vec{f}_{2}^{0}=[0,1,0]\,,\qquad\qquad\vec{f}_{2}^{-}=[0,0,-1]\,.$ (291)
@@ 7.2 Ring relations
Simple products such as $[L_{1}^{+}]^{a}[L_{2}^{+}]^{b}[L_{3}^{+}]^{c}$, $[L_{1}^{+}]^{a}[L_{2}^{+}]^{b}[L_{3}^{0}]^{c}$, $[L_{1}^{+}]^{a}[L_{2}^{0}]^{b}[L_{1}^{-}]^{c}$, etc. also give linearly-independent candidates framed BPS degeneracies of half-BPS line defects with general tropical charge.
@@ 7.4 Tensor products
00 0 $\longrightarrow\mathfrak{q}B_{i+2}^{0\,\prime}$ $\longrightarrow B_{i}^{+}\otimes_{A}B_{i+1}^{-}$ $\longrightarrow A$ $\longrightarrow 0\,,$ (374)
00 0 $\longrightarrow A$ $\longrightarrow B_{i+1}^{-}\otimes_{A}B_{i}^{+}$ $\longrightarrow\mathfrak{q}^{-1}B_{i+2}^{0}$ $\longrightarrow 0\,,$ (375)
00 0 $\longrightarrow A$ $\longrightarrow B_{i+1}^{+}\otimes_{A}B_{i}^{-}$ $\longrightarrow\mathfrak{q}^{-1}X_{-\gamma_{1}-\gamma_{3}}B_{i+2}^{0}$ $\longrightarrow 0\,,$ (376)
00 0 $\longrightarrow\mathfrak{q}X_{-\gamma_{1}-\gamma_{3}}B_{i+2}^{0}$ $\longrightarrow B_{i}^{-}\otimes_{A}B_{i+1}^{+}$ $\longrightarrow A$ $\longrightarrow 0\,,$ (377)
00 0 $\longrightarrow\mathfrak{q}B_{i-2}^{+}$ $\longrightarrow B_{i}^{+}\otimes_{A}B_{i}^{0}$ $\longrightarrow B_{i-1}^{+}$ $\longrightarrow 0\,,$ (378)
00 0 $\longrightarrow B_{i-1}^{+}$ $\longrightarrow B_{i}^{0}\otimes_{A}B_{i}^{+}$ $\longrightarrow\mathfrak{q}^{-1}B_{i-2}^{+}$ $\longrightarrow 0\,,$ (379)
00 0 $\longrightarrow B_{i+1}^{-}$ $\longrightarrow B_{i}^{-}\otimes_{A}B_{i}^{0}$ $\longrightarrow\mathfrak{q}^{-1}B_{i+2}^{-}$ $\longrightarrow 0\,,$ (380)
00 0 $\longrightarrow\mathfrak{q}B_{i+2}^{-}$ $\longrightarrow B_{i}^{0}\otimes_{A}B_{i}^{-}$ $\longrightarrow B_{i+1}^{-}$ $\longrightarrow 0\,,$ (381)
00 0 $\longrightarrow\mathfrak{q}X_{\gamma_{1}+\gamma_{3}}B^{+}_{i+2}\otimes_{A}B^{-}_{i+2}$ $\longrightarrow B^{0}_{i}\otimes_{A}B^{0}_{i+1}$ $\longrightarrow X_{\gamma_{1}+\gamma_{3}}A$ $\longrightarrow 0\,,$ (382)
00 0 $\longrightarrow X_{\gamma_{1}+\gamma_{3}}A$ $\longrightarrow B^{0}_{i+1}\otimes_{A}B^{0}_{i}$ $\longrightarrow\mathfrak{q}^{-1}X_{\gamma_{1}+\gamma_{3}}B^{+}_{i+2}\otimes_{A}B^{-}_{i+2}$ $\longrightarrow 0$ (383)
@@ 7.4.1 Checking $\mathbb{Z}_{3}$ symmetry
    and noting that $u$ has a spin $\frac{1}{2}$, we see that $u$ generates a submodule isomorphic to $\mathfrak{q}B_{2}^{0}$ after twisted by $(-1)^{\langle\gamma_{1},\gamma\rangle}$. [^22] Recall that $\mathfrak{q}$ by itself caries a twist of every generator by $-1$. Once we quotient out the sub-module generated by $u$, $v$ is cyclic and generates $A$ [^23] Or more precisely, the image of $v$. This can be summed up in the short exact sequence
@@
    $\tilde{v}$ commutes with all $e^{(i)}_{n}$’s while $\tilde{u}$ satisfies the same exchange relations (387) as $u$. The vector $\tilde{v}$ generates a submodule isomorphic to $A$ while, in the quotient, $\tilde{u}$ generates $B_{2}^{0}$. Finally, noting that the spin of $u$ is $-\frac{1}{2}$, we conclude
@@
    On the other hand, $u$ commutes with all the $e^{(1)}_{n}$’s except that $[u,e^{(1)}_{n}]=v^{\prime\prime}$, commutes with all $e^{(2)}_{n}$’s, and anti-commutes with all the $e^{(3)}_{n}$’s except that $\{u,e^{(3)}_{0}\}=ve^{(3)}_{0}e^{(2)}_{0}$. Therefore after modding out this submodule generated by $v$, the class of $u$ generates $A$ up to a sign twist, producing
@@
    implying that $u$ generates submodule isomorphic to twisted $A$. What’s left to prove is that in the quotient, the class of $v$ behaves as $b_{4}^{0}$ and that its spin is $-\frac{1}{2}$. One can check that modulo $u\otimes A$, $v$ satisfies the same equation for the $v\equiv b_{2}^{+}\otimes b_{3}^{-}$ for the opposite product (403):
@@
$[p(s_{\bullet,\bullet})]_{\vec{n}}\to[p(s_{\bullet,\bullet}-z)]_{\vec{n}}\,.$ (413)
@@
We can also define bimodules $\mathrm{Sh}_{0,\mathbb{C}}(Q;\vec{f},\vec{f}^{\prime})$ as a modification of $\mathrm{Sh}_{\mathbb{C}}(Q;\vec{f}+\vec{f}^{\prime})$ where we use $s_{i,a}$ for $\vec{f}$ factors and $(s_{i,a}-z)$ for $\vec{f}^{\prime}$ factors in the shuffle product. Intuitively, this represents bound states to two heavy particles, placed at 00 and $z$ in $\mathbb{C}$. Then $\mathrm{sH}_{0,\mathbb{C}}(Q;\vec{f},\vec{f}^{\prime})$ is defined in the same manner. It is the natural IR description of two line defects separated in the holomorphic direction.
@@ A.1 Framed BPS degeneracies via quiver Yangian algorithm
These $e$’s are fermionic since $|i|=1$. The $e-e$ mode relations from (419) is then
@@
[…]
@@
ketΠi
[…]
In this notation, each $i$Box is a path from the framing node * to a vertex $i\in Q_{0}$, and its equivariant weight is the sum of the weights of all arrows along the corresponding path h(i
@@ A.1.3 Example: $\hat{A}_{2}$ quiver
$\varphi^{1\Leftarrow 1}(z)=1\,,\qquad\qquad\,\,\,$ $\varphi^{1\Leftarrow 2}(z)=z+h_{1}\,,\qquad\,\varphi^{1\Leftarrow 3}(z)=\frac{1}{z-h_{3}}\,,$ (450)
$\varphi^{2\Leftarrow 1}(z)=\frac{1}{z-h_{1}}\,,\qquad$ $\varphi^{2\Leftarrow 2}(z)=1\,,\qquad\qquad\,\,\,\varphi^{2\Leftarrow 3}(z)=z+h_{2}\,,$ (451)
$\varphi^{3\Leftarrow 1}(z)=z+h_{3}\,,\qquad\,$ $\varphi^{3\Leftarrow 3}(z)=\frac{1}{z-h_{2}}\,,\qquad\varphi^{3\Leftarrow 3}(z)=1\,,$ (452)
@@
    where the zero at $z=h_{2}$ in $\psi^{(2)}_{0}(z)=\varphi^{2\Leftarrow*}(z)$ is due to the loop 12$\rightarrow$ 1$\rightarrow$ 2$\rightarrow$*∗ → 1 → 2 → ∗. This gives

### v3

@@ 5.1 Framed quivers and BPS degeneracies
[…]
@@ 6.2 Simple products
The auxiliary bimodule map $\pi_{\vec{f}_{a},\vec{f}_{a+1}}$ that we defined above indeed maps $b\otimes b\to b$. Then we still need to show that $\pi_{\vec{f}_{a},\vec{f}_{a+1}}$ actually lands in $\mathrm{sH}(Q;\vec{f}_{a}+\vec{f}_{a+1})$ rather than $\mathrm{Sh}(Q;\vec{f}_{a}+\vec{f}_{a+1})$, i.e. that we do not encounter images of the form babbab where $a$ is a non-trivial algebra element that cannot be brought to the left and/or to the right of the $b$’s.
@@ 6.3 Interesting tensor products
    $e^{(1)}_{n}\,,\,b_{\pm\gamma_{1}}]=0\qquad\textrm{and}\qquad e^{(2)}_{n}\,b_{\pm\gamma_{1}}=b_{\pm\gamma_{1}}\,e^{(2)}_{n\mp 1}\,$. (239)
@@ 6.3.3 $B_{2}$ and $B_{4}$
anti-commutes with all generators in AA. [^22] If we had done this computation directly in terms of modes using only the exchange relations, we would have $\{e^{(1)}_{n\geq 1},u_{24}\}=\{e^{(2)}_{n},u_{24}\}=0$, and $\{e^{(1)}_{0},u_{24}\}=-b_{\gamma_{1}}\otimes c_{0}$, which vanishes only upon invoking the cubic relation (263). Therefore we get the expected
@@ 6.3.5 $B_{4}$ and $B_{1}$
anti-commutes with all $e^{(1)}_{n}$’s and commutes with all $e^{(2)}_{n}$’s. [^23] Similar to the $B_{2}\otimes_{A}B_{4}$ case, if we had done this computation directly in terms of modes using only the exchange relations, we would have $\{e^{(1)}_{n},u_{41}\}=[e^{(2)}_{n\geq 1},u_{41}]=0$, and $e^{(2)}_{0},u_{41}]=-c_{0}\otimes b_{\gamma_{2}}$, which vanishes only upon invoking the cubic relation (263). Therefore we get the exact sequences
@@ 7.1 Mutation orbits
$1,0,1]\quad\rightarrow\quad[1,-1,0]\quad\rightarrow\quad[0,1,-1]\,.$ (285)
@@
$\vec{f}_{0}^{+}=[1,0,0]\,$$\qquad\quad$ $\vec{f}_{0}^{0}=[-1,0,1]\,,\qquad\quad\vec{f}_{0}^{-}=[0,0,-1]\,$, (286)
$\vec{f}_{1}^{+}=[0,1,0]\,$$\qquad\quad$ $\vec{f}_{1}^{0}=[1,-1,0]\,,\qquad\quad\vec{f}_{1}^{-}=[-1,0,0]\,$, (287)
$\vec{f}_{2}^{+}=[0,0,1]\,$$\qquad\quad$ $\vec{f}_{2}^{0}=[0,1,-1]\,,\qquad\quad\vec{f}_{2}^{-}=[0,-1,0]\,$, (288)
@@
$\vec{f}_{0}^{+}=[1,0,0]\,$$\qquad\quad$ $\vec{f}_{0}^{0}=[0,-1,0]\,,\,\,\,\,\qquad\,\,\,\vec{f}_{0}^{-}=[0,1,-1]\,$, (289)
$\vec{f}_{1}^{+}=[0,0,1]\,$$\qquad\quad$ $\vec{f}_{1}^{0}=[1,0,-1]\,,\qquad\quad\,\vec{f}_{1}^{-}=[-1,0,0]\,$, (290)
$\vec{f}_{2}^{+}=[1,-1,0]\,$$\qquad\quad\,$ $\vec{f}_{2}^{0}=[0,1,0]\,,\qquad\qquad\vec{f}_{2}^{-}=[0,0,-1]\,.$ (291)
@@ 7.2 Ring relations
Simple products such as $L_{1}^{+}]^{a}[L_{2}^{+}]^{b}[L_{3}^{+}]^{c}$, $L_{1}^{+}]^{a}[L_{2}^{+}]^{b}[L_{3}^{0}]^{c}$, $L_{1}^{+}]^{a}[L_{2}^{0}]^{b}[L_{1}^{-}]^{c}$, etc. also give linearly-independent candidates framed BPS degeneracies of half-BPS line defects with general tropical charge.
@@ 7.4 Tensor products
$0$ $\longrightarrow\mathfrak{q}B_{i+2}^{0\,\prime}$ $\longrightarrow B_{i}^{+}\otimes_{A}B_{i+1}^{-}$ $\longrightarrow A$ $\longrightarrow 0\,$, (374)
$0$ $\longrightarrow A$ $\longrightarrow B_{i+1}^{-}\otimes_{A}B_{i}^{+}$ $\longrightarrow\mathfrak{q}^{-1}B_{i+2}^{0}$ $\longrightarrow 0\,$, (375)
$0$ $\longrightarrow A$ $\longrightarrow B_{i+1}^{+}\otimes_{A}B_{i}^{-}$ $\longrightarrow\mathfrak{q}^{-1}X_{-\gamma_{1}-\gamma_{3}}B_{i+2}^{0}$ $\longrightarrow 0\,$, (376)
$0$ $\longrightarrow\mathfrak{q}X_{-\gamma_{1}-\gamma_{3}}B_{i+2}^{0}$ $\longrightarrow B_{i}^{-}\otimes_{A}B_{i+1}^{+}$ $\longrightarrow A$ $\longrightarrow 0\,$, (377)
$0$ $\longrightarrow\mathfrak{q}B_{i-2}^{+}$ $\longrightarrow B_{i}^{+}\otimes_{A}B_{i}^{0}$ $\longrightarrow B_{i-1}^{+}$ $\longrightarrow 0\,$, (378)
$0$ $\longrightarrow B_{i-1}^{+}$ $\longrightarrow B_{i}^{0}\otimes_{A}B_{i}^{+}$ $\longrightarrow\mathfrak{q}^{-1}B_{i-2}^{+}$ $\longrightarrow 0\,$, (379)
$0$ $\longrightarrow B_{i+1}^{-}$ $\longrightarrow B_{i}^{-}\otimes_{A}B_{i}^{0}$ $\longrightarrow\mathfrak{q}^{-1}B_{i+2}^{-}$ $\longrightarrow 0\,$, (380)
$0$ $\longrightarrow\mathfrak{q}B_{i+2}^{-}$ $\longrightarrow B_{i}^{0}\otimes_{A}B_{i}^{-}$ $\longrightarrow B_{i+1}^{-}$ $\longrightarrow 0\,$, (381)
$0$ $\longrightarrow\mathfrak{q}X_{\gamma_{1}+\gamma_{3}}B^{+}_{i+2}\otimes_{A}B^{-}_{i+2}$ $\longrightarrow B^{0}_{i}\otimes_{A}B^{0}_{i+1}$ $\longrightarrow X_{\gamma_{1}+\gamma_{3}}A$ $\longrightarrow 0\,$, (382)
$0$ $\longrightarrow X_{\gamma_{1}+\gamma_{3}}A$ $\longrightarrow B^{0}_{i+1}\otimes_{A}B^{0}_{i}$ $\longrightarrow\mathfrak{q}^{-1}X_{\gamma_{1}+\gamma_{3}}B^{+}_{i+2}\otimes_{A}B^{-}_{i+2}$ $\longrightarrow 0$ (383)
@@ 7.4.1 Checking $\mathbb{Z}_{3}$ symmetry
    and noting that $u$ has a spin $\frac{1}{2}$, we see that $u$ generates a submodule isomorphic to $\mathfrak{q}B_{2}^{0}$ after twisted by $(-1)^{\langle\gamma_{1},\gamma\rangle}$. [^24] Recall that $\mathfrak{q}$ by itself caries a twist of every generator by 1-1. Once we quotient out the sub-module generated by $u$, $v$ is cyclic and generates AA. [^25] Or more precisely, the image of $v$. This can be summed up in the short exact sequence
@@
    $\tilde{v}$ commutes with all $e^{(i)}_{n}$’s while $\tilde{u}$ satisfies the same exchange relations (387) as $u$. The vector $\tilde{v}$ generates a submodule isomorphic to $A$ while, in the quotient, $\tilde{u}$ generates $B_{2}^{0}$. Finally, noting that the spin of $u$ is $\frac{1}{2}$, we conclude
@@
    On the other hand, $u$ commutes with all the $e^{(1)}_{n}$’s except that $u,e^{(1)}_{n}]=v^{\prime\prime}$, commutes with all $e^{(2)}_{n}$’s, and anti-commutes with all the $e^{(3)}_{n}$’s except that $\{u,e^{(3)}_{0}\}=ve^{(3)}_{0}e^{(2)}_{0}$. Therefore after modding out this submodule generated by $v$, the class of $u$ generates $A$ up to a sign twist, producing
@@
    implying that $u$ generates submodule isomorphic to twisted $A$. What’s left to prove is that in the quotient, the class of $v$ behaves as $b_{4}^{0}$ and that its spin is $\frac{1}{2}$. One can check that modulo $u\otimes A$, $v$ satisfies the same equation for the $v\equiv b_{2}^{+}\otimes b_{3}^{-}$ for the opposite product (403):
@@
$p(s_{\bullet,\bullet})]_{\vec{n}}\to[p(s_{\bullet,\bullet}-z)]_{\vec{n}}\,.$ (413)
@@
We can also define bimodules $\mathrm{Sh}_{0,\mathbb{C}}(Q;\vec{f},\vec{f}^{\prime})$ as a modification of $\mathrm{Sh}_{\mathbb{C}}(Q;\vec{f}+\vec{f}^{\prime})$ where we use $s_{i,a}$ for $\vec{f}$ factors and $(s_{i,a}-z)$ for $\vec{f}^{\prime}$ factors in the shuffle product. Intuitively, this represents bound states to two heavy particles, placed at 0 and $z$ in $\mathbb{C}$. Then $\mathrm{sH}_{0,\mathbb{C}}(Q;\vec{f},\vec{f}^{\prime})$ is defined in the same manner. It is the natural IR description of two line defects separated in the holomorphic direction.
@@ A.1 Framed BPS degeneracies via quiver Yangian algorithm
These $e$’s are fermionic since $|i|=1$. The eee-e mode relations from (419) is then
@@
reproducing the exchange relations of the CoHA (49) upon setting all the equivariant weights $h_{I}=0$. [^28] Note that in general it is important to turn off $h_{I}$ only at the last step, in particular, after moving the denominator of $\varphi^{i\Leftarrow j}(u)$ to the other side of the equation; however, for the 2-acyclic quivers, which are the focus of this paper, this subtlety is absent since there the bonding factors $\varphi^{i\Leftarrow j}$ have either only numerators or only denominators. (The other situation $\langle\gamma_{i},\gamma_{j}\rangle=-|j\rightarrow i|\leq 0$ can be treated in the same way.) Therefore, the subalgebra of the quiver Yangian that is generated by the $e$’s can be viewed as the equivariant version of the spherical CoHA. [^29] Conversely, for quivers (with potentials) that arise from IIA on certain toric Calabi-Yau threefolds without compact four cycles — not the type that we are considering in this paper — the Drinfeld double of the equivariant spherical CoHA (after quotienting out certain relations on the Cartan generators) was shown to be isomorphic to the corresponding quiver Yangian Rapcak:2020ueh .
@@
In this notation, each $i$ $\Box$ is a path from the framing node * to a vertex $i\in Q_{0}$, and its equivariant weight is the sum of the weights of all arrows along the corresponding path h(i
@@ A.1.3 Example: $\hat{A}_{2}$ quiver
$\varphi^{1\Leftarrow 1}(z)=1\,$$\qquad\qquad\,\,\,$ $\varphi^{1\Leftarrow 2}(z)=z+h_{1}\,,\qquad\,\varphi^{1\Leftarrow 3}(z)=\frac{1}{z-h_{3}}\,$, (450)
$\varphi^{2\Leftarrow 1}(z)=\frac{1}{z-h_{1}}\,$$\qquad$ $\varphi^{2\Leftarrow 2}(z)=1\,,\qquad\qquad\,\,\,\varphi^{2\Leftarrow 3}(z)=z+h_{2}\,$, (451)
$\varphi^{3\Leftarrow 1}(z)=z+h_{3}\,$$\qquad\,$ $\varphi^{3\Leftarrow 3}(z)=\frac{1}{z-h_{2}}\,,\qquad\varphi^{3\Leftarrow 3}(z)=1\,$, (452)
@@
    where the zero at $z=h_{2}$ in $\psi^{(2)}_{0}(z)=\varphi^{2\Leftarrow*}(z)$ is due to the loop $\rightarrow 1\rightarrow 2\rightarrow*$. This gives